If a line with direction ratios 2:2:1 intersects the line x-7/2=y-5/2=z-3/1 and x-1/2=y+1/4=z+1/3 at A and B then AB =

Grade:12

1 Answers

Deepak Kumar Shringi
6 years ago
While posting questions please once confirm that you are posting a coorect question here in this question as you written dr with 2,2,1 intersecting line another line you given with dr 2,2,1 that are parellel to each other i mean (x-7/2=y-5/2=z-3/1) so intersection is not possible

Thanks